OpenStudy (anonymous):

Find the percent of decrease. Round to the nearest tenth, if necessary. from 23 cm to 17 cm a. 26.1% decrease c. 383.3% decrease b. 135.3% decrease d. 35.3% decrease

OpenStudy (anonymous):

\[\frac{change}{original} = \% change\] \[\frac{23-17}{23} = ?\%\]

OpenStudy (anonymous):

what?